# The earliest phases of CNO enrichment in galaxiesthanks: Full Tables 1 and 2 are only available at the CDS via anonymous ftp to cdsarc.u-strasbg.fr(130.79.128.5) or via http://cdsarc.u-strasbg.fr/viz-bin/cat/J/A+A/xxx/Ayyy

### v1

@@ 8 Conclusions
To explore the earliest phases of chemical enrichment in galaxies, we implement a stochastic star-formation component and detailed Pop III nucleosynthesis prescriptions in a GCE model that already reproduces satisfactorily well the main observed properties of our Galaxy (see Sects. 3.1 and 4). By comparing our model predictions with a homogeneous stellar sample of unmixed stars (Sect.2.1.1), whose chemical abundances reflect the composition of their natal birth clouds, we first demonstrate that the scatter in CNO abundances and abundance ratios observed in the MW halo can be largely attributed to the contributions of Pop III SNe with different masses and explosion energies. Then, we extend the model to the high-redshift universe and predict the chemical abundance ratios for a proto-MW galaxy seen at different redshifts (Sect. 6.1). Finally, we adapt the model and provide an alternative interpretation of the chemical abundances observed in GN-z11 (Sect. 6.2). Our key findings can be summarized as follows:
@@ Appendix A Predictions for MW analogs at different redshifts
In Fig.10 we display the density maps in different plane $\rm log(N/O)$, $\rm log(C/O)$ and $\rm log(C/N)$ versus $\rm log(O/H)+12$ for a Milky Way-like galaxy at different redshift, up to $z\sim 2$. Following the initial chemical enrichment by Pop III SNe, which establishes a wide range of initial abundances, a phase of rapid chemical enrichment occurs. The density contours indicate that most clumps merging to form the halo achieve values of approximately $\rm log(N/O)\sim-1$, $\rm log(C/O)\sim-0.5$, $\rm log(C/N)\sim 0.5$ at $\rm log(O/H)+12\sim 8.5$ after only about 400 Myr (at $z\sim 11$). The subsequent evolution proceed slowly. At $z\sim 2$ the gas reaches $\rm\log(N/O)\sim-0.5$, $\rm\log(C/O)\sim-0.1$, $\rm\log(C/N)\sim 0.5$ and $\rm\log(O/H)+12\sim 9.1$ reflecting the contribution of subsequent stellar generations.

### v2

@@ 7.2 C- and N-enhanced high-redshift systems
As for the class of the rarer extreme C-emitters (see Schaerer et al., 2024, their figure 2), we study the emblematic case of GS-z12. Also in this case, massive Pop III stars that explode as faint SNe are found to play a key role in reproducing the observed high C/O ratio at low metallicity, notwithstanding the lowest stellar mass reached by GS-z12.
@@ 8 Conclusions
To explore the earliest phases of chemical enrichment in galaxies, we implement a stochastic star-formation component and detailed Pop III nucleosynthesis prescriptions in a GCE model that already reproduces satisfactorily well the main observed properties of our Galaxy (see Sects. 3.1 and 4). By comparing our model predictions with a homogeneous stellar sample of unmixed stars (Sect.2.1.1), whose chemical abundances reflect the composition of their natal birth clouds, we first demonstrate that the scatter in CNO abundances and abundance ratios observed in the MW halo can be largely attributed to the contributions of Pop III SNe with different masses and explosion energies. Then, we extend the model to the high-redshift universe and predict the chemical abundance ratios for a proto-MW galaxy seen at different redshifts (Sect. 6.1). Finally, we adapt the model and provide an alternative interpretation of the chemical abundances observed in GN-z11 and GS-z12 (Sect. 6.2). Our key findings can be summarized as follows:
@@
  • (vi)
    Pre-enrichment by faint Pop III SNe could also be the key to explain the high C/O ratios observed in young low-metallicity systems, such as GS-z12.
@@
9 Data availability
Tables 1 and 2 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Appendix A Predictions for MW analogs at different redshifts
In Fig. 13 we display the density maps in different planes, $\rm log(N/O)$, $\rm log(C/O)$, and $\rm log(C/N)$ versus $\rm log(O/H)+12$, for a Milky Way-like galaxy at different redshifts, up to $z\sim 2$. Following the initial chemical enrichment by Pop III SNe, which establishes a wide range of initial abundances, a phase of rapid chemical enrichment occurs. The density contours indicate that most clumps merging to form the halo achieve values of approximately $\rm log(N/O)\sim-1$, $\rm log(C/O)\sim-0.5$, $\rm log(C/N)\sim 0.5$ at $\rm log(O/H)+12\sim 8.5$ after only about 400 Myr (at $z\sim 11$). The subsequent evolution proceeds slowly. At $z\sim 2$ the gas reaches $\rm\log(N/O)\sim-0.5$, $\rm\log(C/O)\sim-0.1$, $\rm\log(C/N)\sim 0.5$ and $\rm\log(O/H)+12\sim 9.1$ reflecting the contribution of subsequent stellar generations.